If you wish to accumulate $205,000 in 6 years, how much must you deposit today in an account that pays a quoted annual interest rate of 15% with monthly compounding of interest?

Question 13 options:

$76,353.67

$88,627.16

$117,874.12

$83,813.03

$147,121.08
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

205000=P*(1+0.15/12)^(12*6)

P=205000/(1+0.15/12)^(12*6)

=205000*0.408844071

=$83813.03(Approx).